OpenStudy (anonymous):

AP Statistics Help Will give Medal! Margin of Error/ Confidence Intervals: For a single proportion the margin of error of a confidence interval is largest for any given sample size n and confidence level C when p-hat = 0.5. This led us to use p-hat = 0.5 for planning purposes. Use these conservative values in the following calculations, and assume that the sample sizes of n. Calculate the margins of error of the 99% confidence intervals for the following choices of n: 10, 30, 50, 100, 200, and 500. Present the results in a table. Summarize your conclusions.

OpenStudy (anonymous):

margin of error = critical value * SD. For a proportion, SD is given by \[\sigma = \sqrt{\frac{\hat{p}(1-\hat{p})}{n}}\] Your critical value is where \(P(Z < z^*) = \frac{1 + 99\%}{2}\).

OpenStudy (anonymous):

Thank you very much! I This helps a lot

OpenStudy (anonymous):

I just didn't know where things were supposed to go :)